OpenStudy (anonymous):

Divide five thousand, nine hundred thirty-four divided by ten to the fourth power..

OpenStudy (danjs):

\[\frac{ 5934 }{ 10^4 } = 5934*10^{-4}\]

OpenStudy (danjs):

negative power, move the decimal to the left